15 litres of oxygen is ozonated. Final volume of mixture is 12 litres. Find percentage yield of reaction.

Grade:11

2 Answers

ULFAT
8 years ago
p { margin-bottom: 0.25cm; line-height: 120%; }

GIVEN OXYGEN = 15L

FINAL VOLUME OF MIXTURE=12L

 

O2 + [O] →O3

1 MOLE OF O2 → 1 MOLE OF O3

16×2 16×3

32 → 48

15L OF O2 → X

15L=15/1000=0.015

32 → 48

LITRES CHANGES INTO GRAMS

0.015→X

X=0.015×48÷3X =0.0225

DENSITY OF O2 = 1.429

V= 12÷1000 = 0.012×1.429

=0.017148gms

v÷x = 0.017148÷0.0225

0.07621333×100 = 76.2133

PERCENTAGE YIELD REACTION= 76.2133

 

DEEPTHI JANGA
8 years ago
 

GIVEN OXYGEN = 15L

FINAL VOLUME OF MIXTURE=12L

 

O2 + [O] →O3

1 MOLE OF O2 → 1 MOLE OF O3

16×2 16×3

32 → 48

15L OF O2 → X

15L=15/1000=0.015

32 → 48

LITRES CHANGES INTO GRAMS

0.015→X

X=0.015×48÷3X =0.0225

DENSITY OF O2 = 1.429

V= 12÷1000 = 0.012×1.429

=0.017148gms

v÷x = 0.017148÷0.0225

0.07621333×100 = 76.2133

PERCENTAGE YIELD REACTION= 76.2133